# How Good Are We? Evaluating Cell AI Foundation Models in Kidney Pathology with Human-in-the-Loop Enrichment

### v1

@@ 3.3.2 Class-wise Weighted Oversampling
In this work, we addressed the imbalance between human-annotated image patches and pseudo labels from foundation models’ “good” predictions (with a ratio of about 1:50) by applying oversampling. Specifically, we implemented a customized class-based weighted oversampling method, as used in [19]. For a training dataset of $N_{\text{Train}}$ samples, the sampling weight for each training sample $i$ was calculated based on its curated class (“good” or “bad” image patches), as outlined below:
$\begin{split}w_{\text{Class}}(i,\gamma_{s})=\frac{N_{\text{Train}}}{\gamma_{s}\left(\sum_{j\in[1,N_{\text{Train}}]|\ell_{j}=\ell_{i}}1\right)+(1-\gamma_{s})N_{\text{Train}}}\end{split}$ (1)
As denoted, $w_{\text{Class}}(i,\gamma_{s})$ represents the sampling weight assigned to image patch $i$ based on its curated class. Higher weights will be assigned to underrepresented classes, ensuring they are sampled more frequently during training. $\gamma_{s}\in[0,1]$ is the oversampling parameter that controls the strength of oversampling. When $\gamma_{s}=0$, there is no oversampling, and all patches are sampled uniformly. When $\gamma_{s}=1$, the maximum balancing is applied, and patches from underrepresented classes are given the highest weights. In this work, we followed the default setting of $\gamma_{s}=0.85$ in [19]. $\sum_{j\in[1,N_{\text{Train}}]|\ell_{j}=\ell_{i}}1$, this sum counts the number of patches in the training dataset that belong to the same class as patch $i$. Specifically, $\ell_{j}$ and $\ell_{i}$ represent the class labels of patches $j$ and $i$, respectively. The numerator $N_{\text{Train}}$ ensures that the weights are normalized based on the total number of training patches. The expression $\gamma_{s}\left(\sum_{j\in[1,N_{\text{Train}}]|\ell_{j}=\ell_{i}}1\right)+(1-\gamma_{s})N_{\text{Train}}$ controls the balance between classes. When $\gamma_{s}$ is large, equation (1) gives more weight to underrepresented classes (with fewer samples), while for smaller $\gamma_{s}$, equation(1) reduces the effect of class imbalance, leading to more uniform sampling. In summary, this oversampling approach ensures that human annotated samples, such as those with challenging nuclei or lower image quality, are sampled more frequently during training, which helps improve the model’s ability to generalize across all cell nuclei in the kidney.
4 Experiments
4.1 Dataset
Evaluation Dataset. As described in the previous section, the evaluation dataset in this study is highly diverse, comprising 2,542 whole-slide images (WSIs) from both public and private sources. It includes samples from both human and rodent tissue, stained with H&E, PASM, and PAS. After discarding contaminated or poor-quality patches, we randomly extracted four 512 $\times$ 512-pixel patches from each WSI at 40$\times$ magnification, resulting in an evaluation dataset of 8,818 image patches. As illustrated in Fig.1b, 57% (1,449 WSIs) were collected from publicly available sources, including the Kidney Precision Medicine Project (KPMP)[43], NEPTUNE [44], and HUBMAP [45, 46], while the remaining 43% (1,093 WSIs) were acquired from an internal collection at Vanderbilt University Medical Center. To increase the diversity of our dataset, we incorporated WSIs from both human and rodent samples.
@@ 4.1 Dataset
Fine-tuning Dataset. Building on the evaluation dataset construction described above, three foundation models were used for inference and curation on each 512 $\times$ 512 image patch. The “good” prediction masks from all three foundation models, along with their corresponding images, were retrieved as image-pseudo label pairs and added to the fine-tuning dataset. Additionally, a small set of 198 images, rated as “bad” by all foundation models, were manually annotated by pathologists and included in the fine-tuning dataset. This resulted in a total of 12,005 image-label pairs. For the training and validation split, 11,807 pairs were used for training, while a diverse set of 100 pairs was reserved for validation. Lastly, an additional 185 images, sampled from all kidney WSI sources, were annotated by pathologists and designated as the hold-out testing set to evaluate the performance of the fine-tuned foundation models. Each 512 $\times$ 512 image patch contains 50 to 300 cell nuclei, providing substantial signals for effective fine-tuning.
4.2 Experimental Design
@@ 4.2.1 Evaluation of Foundation Model Performance
The segmentation of nuclei instances from the three foundation models was evaluated as “good,” “medium,” or “bad,” following a standard established by a renal pathologist with more than 20 years of experience. We applied this rating system to quantitatively assess each cell foundation model’s predictions in our evaluation dataset. The specific definitions are as follows:
  • Good: Predictions capture approximately 90% of the nuclei in an image patch.
  • Medium: Predictions capture between 50% and 90% of the nuclei in an image patch.
  • Bad: Predictions capture less than 50% of the nuclei in an image patch.
@@
Fused Model Performance Evaluation. Building on the evaluation of each individual cell foundation model’s performance, we conducted a joint model performance evaluation. The rating assignment for each image patch was determined by the fusion of ratings from the three foundation models. Specifically, an image patch was assigned to the “good” prediction class if any model rated it as “good.” Conversely, an image patch was categorized as “bad” only if all foundation models were assigned a “bad” rating. The remaining image patches were categorized as “medium.” Similarly, the distribution of rating classes was analyzed. The results of this analysis indicate the upper bound of applying multiple foundation models to our domain task and highlight the potential for fine-tuning these models through our data enrichment strategies. Furthermore, a taxonomy of common errors made by all cell foundation models in this study can be derived from the fused “bad” image patches.
Cross-Model Performance Evaluation. In this work, we also evaluated cross-model performance on our kidney nuclei dataset by conducting an agreement analysis among the foundation models. The steps are as follows:
  1. 1.
    Agreement Matrix. We computed the percentage of agreement between each pair of models to assess how often they assigned the same rating to image patch predictions. This matrix highlights the consistency among the foundation models in their rating assignments.
  2. 2.
    Class-Specific Agreement. Image patches were categorized into three groups: All Three Models Agree, where all models assigned the same rating; Two Models Agree, where any two models assigned matching ratings; and No Agreement, where none of the models assigned the same rating to the image patch.
@@ 4.2.2 Data-Enriched Fine-Tuning with Multiple Foundation Models
To improve kidney nuclei segmentation performance, we leveraged data enrichment from multiple foundation models’ curation outcomes and fine-tuned the models under three different settings. Each foundation model was fine-tuned using either “good” image patches (foundation model-generated pseudo labels), “bad” image patches (pathologist-corrected), or both. As previously explained, fine-tuning on corrected “bad” image patches aims to narrow the knowledge gap between domains by focusing on samples where all foundation models fail. “Good” image patches enable mutual knowledge distillation across foundation models, enhancing the training of each individual model. The details of the fine-tuning experiments are shown in the Table. 1. For each fine-tuning setting, we used different scales of the training dataset (25%, 50%, 75%, and 100%) to assess the consistency of the model’s fine-tuning. Specifically, each entry in the “Incremental Training Dataset Settings” represents the size of the scaled training dataset used for fine-tuning experiments. For example, when fine-tuning with only “good” image patches derived from foundation models’ predictions, we conducted experiments using 2,952 (25%), 5,901 (50%), 8,854 (75%), and 11,807 (100%) “good” predictions. Similar experiments were performed for all three fine-tuning strategies across the three foundation models: Cellpose, StarDist, and CellViT.
@@
“good” patches only Models’ Predictions 2,952 5,901 8,854 11,807
“bad” patches only Pathologists’ correction 50 99 149 198
“good” + “bad” patches Combine Both 3,002 6,000 9,003 12,005
@@
Table 1: The table summarizes the model fine-tuning experiments conducted under three settings: using “good” image patches (foundation model-generated pseudo labels), “bad” image patches (pathologist-corrected), and a combination of both. Each data-enriched fine-tuning experiment was performed at different scales of the training dataset (25%, 50%, 75%, and 100%) for the three foundation models: Cellpose, StarDist, and CellViT. The numerical values indicate the size of the scaled training dataset for each individual fine-tuning experiment.
4.3 Evaluation Metrics
@@
To assess the instance segmentation performance of the foundation models after fine-tuning, we used Recall (equation(2)), Precision (equation(3)), and F1 score (equation(4)) as evaluation metrics.
Recall measures the ability of the model to correctly identify all nuclei instances in an image patch.
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}=\frac{TP}{TP+FN}$ (2)
Precision measures the proportion of correctly identified nuclei instances among all the instances predicted by the model.
Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}=\frac{TP}{TP+FP}$ (3)
F1 score is calculated as the harmonic mean of precision and recall, providing a single metric that balances both by weighting their average.
$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\times TP}{2\times TP+FP+FN}$ (4)
Moreover, the Intersection over Union (IoU) threshold is set to 0.5 to consider a detection as a True Positive (TP). $|TP|,|FP|,\text{and }|FN|$ are the number of True Positives, False Positives, and False Negatives, respectively.
  • True Positives (TP): Matched pairs of nuclei, i.e., correctly detected nuclei.
  • False Positives (FP): Unmatched predicted nuclei, i.e., predicted nuclei without matching ground truth (GT) nuclei.
  • False Negatives (FN): Unmatched GT nuclei, i.e., GT nuclei without matching predicted nuclei.
4.4 Implementation Details
@@
Inference: For a simple implementation of model inference using each foundation model’s released pretrained weights with GPU support, our previous work [51] provides customized object-oriented python modules. Models were trained using Python 3.9 and PyTorch 2.0.1, with GPU acceleration provided by CUDA 11.7. The experiments were conducted on an NVIDIA RTX A5000 GPU with 24GB of memory.
@@
To maintain consistency in experimental settings and enable better comparison of fine-tuned performance, we applied the same oversampling factor $\gamma_{s}=0.85$ across all experiments that used both “good” and “bad” image patches. Due to the imbalance between “good” and “bad” image patches, the number of training epochs was set to 50 for all fine-tuning experiments involving “good” image patches, while experiments using only “bad” image patches were trained for 25 epochs. The batch size was consistently set to 16 across all experiments.
Cellpose Training: For training Cellpose using our true color images, we first converted the images into the format required for training the Cellpose nucleus model. Each image was converted into a 2-band format, where the first band represents the grayscale image and the second band consists entirely of zeros. This follows the default Cellpose nucleus model inference format, where the first channel is used for nuclei segmentation. Due to different loss convergence patterns in training Cellpose, the base learning rate was set to 0.1 for fine-tuning with only “bad” image patches and 0.001 for fine-tuning that includes “good” image patches. Other experimental settings are unchanged according to  [17].
StarDist Training: The learning rate was maintained at 0.0003 (default), with all other settings unchanged as in [18]. In this work, the fine-tuned StarDist model can be converted and loaded into QuPath extension format for easy validation and labeling purposes.
CellViT Training: The base learning rate was set to 0.0003, with a scheduling factor of 0.85 to gradually reduce the learning rate during training. HoVer-Net post-processing with equal weighting on each loss objective (excluding nuclei classification and tissue classification) was applied during training. The pretrained ViT256 encoder and decoder were trained end-to-end, with no layers frozen. All other hyperparameters remained the same as in the original setting [19].
@@ 5.1 Evaluation of Foundation Model Performance
Individual Model Performance Evaluation. First, we present the evaluation of each individual foundation model’s performance. We analyzed the distribution of human-rated predictions for each foundation model across our evaluation dataset. From the barplot in the upper panel of Fig. 3, we can observe the distribution of the rated predictions for each foundation model in this study. For all three foundation models (Cellpose, StarDist, and CellViT), at least 40% of predictions are rated as “bad” or “medium.” at least 40% of predictions are “good.” This indicates that, despite not being specifically trained on kidney pathology nuclei data, the models retain some transferable knowledge applicable to our task. However, none of the foundation models is perfect, and the 40% performance gap highlights the ongoing need for more specialized foundation models tailored specifically to kidney pathology and their challenging cell nuclei segmentation cases. Among the evaluated cell foundation models, Cellpose and StarDist have similar amounts of “good” rated predictions, while CellViT shows a notably higher number of “good” predictions of 5,609 (63%), suggesting its potentially superior performance and adaptability in kidney pathology.
Fused Model Performance Evaluation. As shown in Fig. 3, building on the evaluation of each individual cell foundation model’s performance, we conducted a fused model performance evaluation through a data enrichment strategy. The rating assignment for each image patch was determined by the fusion of ratings from the three foundation models. This evaluation aims to demonstrate the potential of applying multiple foundation models together in our domain task. As shown, after data enrichment, the fused model performance indicates an increase in the number of “good” image patches (6,001 “good” image patches, representing 68% of the dataset) and a decrease in the number of “bad” image patches (534 image patches, representing 6% of the dataset). Consequently, the “good” image patches surpass those of each foundation model, while the “medium” and “bad” image patches fall below all three models. This highlights the expanded capacity of the foundation models and supports subsequent continual learning using data-enriched prediction outcomes.
@@
Furthermore, as illustrated in the lower panel of Fig. 3, a taxonomy of common errors made by all cell foundation models is summarized from the fused set of “bad” image patches. It is evident that certain types of nuclei and tissue conditions diminish the segmentation performance of foundation models. Specifically, long and flat nuclei, nuclei with blurred boundaries, and densely distributed nuclei within glomeruli are particularly difficult to segment accurately. Additionally, slides with faint staining, those containing fatty tissues, or slides with an excessive number of red blood cells can result in a higher incidence of model failures.
Cross-Model Performance Agreement. Next, we move beyond evaluating the performance of individual foundation models and assess their collective behavior by examining the agreement and disagreement among their predictions. Fig. 4a shows the consistency between model predictions. As illustrated, no pair of models reaches over 90% agreement or complete disagreement. The highest agreement occurs between Cellpose and StarDist (0.76), while the lowest is between Cellpose and CellViT (0.67). These non-extreme values, ranging from 0.67 to 0.76, suggest that the current SOTA nuclei foundation models in digital pathology do not generalize exceptionally well to large-scale, diverse kidney datasets. Each model exhibits its own strengths, highlighting the potential for combining multiple SOTA foundation models to improve performance in downstream kidney nuclei tasks. Fig. 4b shows the percentages of image patches where all three models agree, two models agree, or no models agree for each rated prediction class (“good”, “medium”, “bad”). It is evident that “bad” samples exhibit the highest inter-model reliability, with over 70% agreement among all three models, indicating the nature of our design that prioritizes human annotation effort for consensus “bad” image patches.
@@
To sum up, by combining the performance evaluations from the results above, we found that none of the foundation models in this study is fully capable of addressing all cell nuclei segmentation challenges in the kidney. This suggests that current foundation model-based segmentation methods are still insufficient to fully solve nuclei segmentation across all organs or cell types in digital histology.
@@
good 25% 0.7297 0.8313 0.6649
good 50% 0.7357 0.8286 0.6748
good 75% 0.7422 0.8447 0.6747
good 100% 0.7529 0.8349 0.6986
bad 25% 0.5353 0.8095 0.4170
bad 50% 0.6444 0.7959 0.5592
bad 75% 0.5671 0.8401 0.4461
bad 100% 0.5807 0.8121 0.4699
good 25% + bad 25% 0.7228 0.8475 0.6466
good 50% + bad 50% 0.7270 0.8481 0.6493
good 75% + bad 75% 0.7352 0.8535 0.6953
good 100% + bad 100% 0.7350 0.8512 0.6601
@@
good 25% 0.7926 0.8989 0.7209
good 50% 0.7931 0.9084 0.7166
good 75% 0.7895 0.9063 0.7110
good 100% 0.7943 0.8991 0.7234
bad 25% 0.8086 0.8862 0.7423
bad 50% 0.8166 0.8834 0.7696
bad 75% 0.8109 0.8864 0.7601
bad 100% 0.8229 0.8847 0.7802
good 25% + bad 25% 0.7980 0.9017 0.7249
good 50% + bad 50% 0.8110 0.9015 0.7464
good 75% + bad 75% 0.8172 0.8961 0.7588
good 100% + bad 100% 0.8182 0.8901 0.7664
@@
good 25% 0.7836 0.8216 0.7621
good 50% 0.7828 0.8296 0.7520
good 75% 0.7854 0.8284 0.7587
good 100% 0.7794 0.8270 0.7476
bad 25% 0.7796 0.7866 0.7857
bad 50% 0.7770 0.7898 0.7791
bad 75% 0.7881 0.7770 0.8137
bad 100% 0.7861 0.7809 0.8070
good 25% + bad 25% 0.7893 0.8259 0.7695
good 50% + bad 50% 0.7895 0.8276 0.7652
good 75% + bad 75% 0.7866 0.8303 0.7580
good 100% + bad 100% 0.7952 0.8302 0.7731
Table 2: Fine-tuning Results for Cellpose, StarDist, and CellViT Models. This table consists of three subtables summarizing the instance segmentation performance metrics (F1 score, Precision, and Recall) for each foundation model across three data enrichment experimental settings (using only “good” image patches, only pathologist-corrected “bad” image patches, and a combination of both). Each subtable compares the baseline performance (highlighted in blue) of the model against its fine-tuned counterparts. The highest values of each evaluation metric for the fine-tuned models are highlighted in red.
Figure 5: Comparison of Fine-Tuning results across three foundation models. Each column represents the fine-tuning of an individual foundation model, and each row compares an instance segmentation performance metric (F1 score, Precision, or Recall) across the models. The X-axis in each plot corresponds to the incremental percentage (25%, 50%, 75%, and 100%) of the training dataset used for each fine-tuning strategy, and the Y-axis shows the performance metrics.
Figure 6: Comparison of Best Fine-Tuned Models Against Baseline Performance. Each foundation model’s baseline is highlighted in pink, while our best data-enriched fine-tuned model is shown in blue. This indicates that all models enhance segmentation performance after fine-tuning. Notably, the baseline model with the highest F1 score before fine-tuning (CellViT) does not yield the best segmentation outcomes after fine-tuning. In contrast, StarDist benefits the most from the proposed data enrichment strategies, achieving the highest F1 score of 0.8229.

### v2

@@ 2.4.2 Class-wise Weighted Oversampling
To mitigate the severe imbalance between “easy” and “hard” patches ($\approx$ 50:1), we adopt the customized weighted-oversampling strategy described in [29]. By controlling an oversampling factor $\gamma_{s}\in[0,1]$, higher sample weights will be assigned to underrepresented class (“hard”), ensuring they are sampled more frequently during training. Thus, challenging nuclei samples are emphasized while catastrophic forgetting is reduced by retaining large-scale “easy” labels. A detailed derivation of the weighted oversampling method is provided in the Supplementary Methods.
2.5 Experimental Setup
2.5.1 Dataset
[…]
@@ 2.5.1 Dataset
Fine-tuning Dataset. Building on the evaluation dataset construction described above, three foundation models were used for inference and curation. “Good” prediction masks and corresponding images from all three foundation models were collected as “easy” image–pseudo-label pairs and added to the fine-tuning dataset. Additionally, a small set of 198 hard image patches, which were consistently rated as “bad” by student raters across all foundation models, were manually annotated by pathologists and included in the fine-tuning dataset. As shown in Table 1, This resulted in a total of 12,005 image-label pairs. For the training and validation split, 11,807 pairs were used for training, while a diverse set of 100 pairs was reserved for validation. Lastly, an additional 185 images, sampled from all kidney WSI sources, were annotated by pathologists and designated as the hold-out testing set to evaluate the performance of the fine-tuned foundation models. Each 512 $\times$ 512 image patch contains 50 to 300 cell nuclei, providing substantial signals for effective fine-tuning. We ensure that no overlap existed between WSIs used for training/validation and those in the hold-out test set.
@@ 2.5.2 Evaluation of Foundation Model Performance
Nuclei instance segmentation results from the three foundation models were evaluated using the criteria shown in Fig. 3. This rating system quantitatively assesses each model’s predictions on the evaluation dataset.
@@
Fused Model Performance Evaluation. Building on the evaluation of each individual cell foundation model’s performance, we conducted a joint model performance evaluation. Specifically, an image patch was assigned to the “good” prediction class if any model rated it as “good.” Conversely, an image patch was categorized as “bad” only if all foundation models were assigned a “bad” rating. The remaining image patches were categorized as “medium.” This joint analysis indicates the upper bound of applying multiple foundation models to our domain-specific task and highlight the potential for fine-tuning these models through our data enrichment strategies. Furthermore, a taxonomy of common errors made by all cell foundation models in this study can be derived from the fused “bad” image patches.
Cross-Model Performance Evaluation. In this work, we also evaluated cross-model performance on our kidney nuclei dataset by conducting an agreement analysis among the foundation models. First, we computed an Agreement Matrix that quantified the percentage of agreement between each pair of models, indicating how often they assigned the same rating to image patch predictions. This matrix highlights the consistency among the foundation models in their rating assignments. Next, we examined Class-Specific Agreement by categorizing image patches into three groups: All Three Models Agree, where all models assigned the same rating; Two Models Agree, where any two models assigned matching ratings; and No Agreement, where none of the models assigned the same rating to the image patch.
@@ 2.5.3 Data-Enriched Fine-Tuning with Multiple Foundation Models
As described earlier, Each foundation model was fine-tuned using either “easy” image patches (foundation model-generated pseudo labels), “hard” image patches (pathologist-corrected), or both. Table 1 shows the details of the fine-tuning experiments. For each fine-tuning setting, we used different scales of the training dataset. Specifically, we selected common proportions—25%, 50%, 75%, and 100%—to evaluate whether the impact of annotation strategies across the three models remains consistent under different dataset sizes.
@@
“Easy” patches only Models’ Predictions 2,952 5,901 8,854 11,807
“Hard” patches only Pathologists’ correction 50 99 149 198
“Easy” + “Hard” patches Combine Both 3,002 6,000 9,003 12,005
@@
Table 1: The table summarizes the model fine-tuning experiments conducted under three settings: using “easy” image patches (foundation model-generated pseudo labels), “hard” image patches (pathologist-corrected), and a combination of both. Each data-enriched fine-tuning experiment was performed at different scales of the training dataset (25%, 50%, 75%, and 100%) for the three foundation models: Cellpose, StarDist, and CellViT. The numerical values indicate the size of the scaled training dataset for each individual fine-tuning experiment.
2.5.4 Evaluation Metrics and Implementation Details
@@
Evaluation metrics. To assess the instance segmentation performance of the foundation models after fine-tuning, we used Recall, Precision, and F1 score as evaluation metrics. Details are provided in the Supplementary Experiments.
@@
Implementations. Both training and inference experiments were implemented using Python 3.9 and PyTorch 2.0.1, with GPU acceleration provided by CUDA 11.7. The experiments were conducted on an NVIDIA RTX A5000 GPU with 24GB of memory. For a simple implementation of model inference using each foundation model’s released pretrained weights with GPU support, our previous work [43] provides customized object-oriented python modules.
@@
During fine-tuning, to ensure consistency in experimental settings and enable meaningful comparison, we applied the same oversampling factor $\gamma_{s}=0.85$ across all experiments that used both “easy” and “hard” patches. Due to the imbalance between “easy” and “hard” samples, the number of training epochs was set to 50 for all fine-tuning experiments involving “easy” image patches, while experiments using only “hard” image patches were trained for 25 epochs. The batch size was consistently set to 16 across all experiments. Detailed information about model fine-tuning is provided in the Supplementary Experiments.
2.6 Statistics and Reproducibility
Model evaluation was performed on 8,818 image patches sampled from 2,542 curated kidney whole slide images. Each patch was assigned a qualitative rating (“good,” “medium,” or “bad”) using established criteria. For inference with the publicly available Cellpose, StarDist, and CellViT models, we used our previously published object-oriented Python modules [43] to ensure consistent preprocessing and execution. To assessed the fine-tuned models, instance segmentation performance was assessed on the test dataset using Recall, Precision, and F1 score, calculated at the image-patch level with mean and standard deviation reported. Training parameters were standardized across models to enable fair comparison. Replicates refer to independent evaluations of each model on the same test dataset under identical inference conditions. All experiments were deterministic, with identical hardware to ensure reproducibility.
2.7 Ethical Consideration
This work was conducted with approval and oversight from the Vanderbilt University Institutional Review Board (IRB 230223). This is a retrospective study on archival tissue. No interventions and experimental manipulations were included. We did not recruit participants. We retrieved the archival kidney tissue biopsy samples from the tissue bank. We used pre-collected and de-identified microscopy data only without knowing the information (e.g., age, genotypic information, past and current diagnosis and treatment categories) of the population characteristics.
@@ 3.1 Evaluation of Foundation Model Performance
Individual Model Performance Evaluation. First, we rated the segmentation performance of each foundation model shown in the upper panel of Fig. 4. CellViT demonstrated superior performance in segmenting nuclei in these kidney images, with 5,609 (63.6%) “good” predictions, which outperforms Cellpose (40.5%) and StarDist (40.2%). This indicates that, despite not being specifically trained on kidney pathology nuclei data, the models retain some transferable knowledge applicable to our task. However, even with the best CellViT model, there is still an existence of approximately 37% of rated predictions from “medium” and “bad,” which underscores the need for a kidney domain-specific foundation model. Thus, the fused evaluation dataset—with enriched “good” image patches (6,001; 68%) and a reduced number of “bad” patches (534; 6%)— were further utilized for continuous fine-tuning.
@@
Taxonomy of Model Failures. As illustrated in the lower panel of Fig. 4, a taxonomy of common errors made by all cell foundation models is summarized from the fused set of “bad” image patches. It is evident long and flat nuclei, nuclei with blurred boundaries, and densely distributed nuclei within glomeruli are particularly difficult to segment accurately. Additionally, slides with faint staining, those containing fatty tissues, or slides with an excessive number of red blood cells can result in a higher incidence of model failures.
Cross-Model Performance Agreement. Next, we assess their collective behavior by examining the agreement and disagreement among their predictions. Fig. 5a shows the consistency between model predictions. As illustrated, no pair of models reaches over 90% agreement or complete disagreement. The highest agreement occurs between Cellpose and StarDist (0.76), while the lowest is between Cellpose and CellViT (0.67). These non-extreme values, ranging from 0.67 to 0.76, suggest that the current SOTA nuclei foundation models in digital pathology do not generalize exceptionally well to large-scale, diverse kidney datasets. Each model exhibits its own strengths, highlighting the potential for combining multiple SOTA foundation models to improve performance in downstream kidney nuclei tasks. Fig. 5b shows the percentages of image patches where all three models agree, two models agree, or no models agree for each rated prediction class (“good”, “medium”, “bad”). It is evident that “bad” samples exhibit the highest inter-model reliability, with over 70% agreement among all three models, indicating the nature of our design that prioritizes human annotation effort for consensus “bad” image patches.
@@
Inter-Rater Agreement. Lastly, we also randomly selected 300 images to assess inter-rater reliability, as shown in Supplementary Table S1. CellViT, which has the strongest baseline performance, achieved the highest agreement of 94.3% between the two raters.
@@
Easy 10% 0.7038 0.8272 0.6276
Easy 25% 0.7297 0.8313 0.6649
Easy 50% 0.7357 0.8286 0.6748
Easy 75% 0.7422 0.8447 0.6747
Easy 100% 0.7529 0.8349 0.6986
Hard 10% 0.5117 0.8236 0.3905
Hard 25% 0.5353 0.8095 0.4170
Hard 50% 0.6444 0.7959 0.5592
Hard 75% 0.5671 0.8401 0.4461
Hard 100% 0.5807 0.8121 0.4699
Easy 10% + Hard 10% 0.7067 0.8444 0.6237
Easy 25% + Hard 25% 0.7228 0.8475 0.6466
Easy 50% + Hard 50% 0.7270 0.8481 0.6493
Easy 75% + Hard 75% 0.7352 0.8535 0.6953
Easy 100% + Hard 100% 0.7350 0.8512 0.6601
@@
Easy 10% 0.7838 0.8962 0.7113
Easy 25% 0.7926 0.8989 0.7209
Easy 50% 0.7931 0.9084 0.7166
Easy 75% 0.7895 0.9063 0.7110
Easy 100% 0.7943 0.8991 0.7234
Hard 10% 0.7876 0.8577 0.7435
Hard 25% 0.8086 0.8862 0.7423
Hard 50% 0.8166 0.8834 0.7696
Hard 75% 0.8109 0.8864 0.7601
Hard 100% 0.8229 0.8847 0.7802
Easy 10% + Hard 10% 0.7899 0.9022 0.7152
Easy 25% + Hard 25% 0.7980 0.9017 0.7249
Easy 50% + Hard 50% 0.8110 0.9015 0.7464
Easy 75% + Hard 75% 0.8172 0.8961 0.7588
Easy 100% + Hard 100% 0.8182 0.8901 0.7664
@@
Easy 10% 0.7687 0.8287 0.7276
Easy 25% 0.7836 0.8216 0.7621
Easy 50% 0.7828 0.8296 0.7520
Easy 75% 0.7854 0.8284 0.7587
Easy 100% 0.7794 0.8270 0.7476
Hard 10% 0.7397 0.7707 0.7295
Hard 25% 0.7796 0.7866 0.7857
Hard 50% 0.7770 0.7898 0.7791
Hard 75% 0.7881 0.7770 0.8137
Hard 100% 0.7861 0.7809 0.8070
Easy 10% + Hard 10% 0.7791 0.8216 0.7498
Easy 25% + Hard 25% 0.7893 0.8259 0.7695
Easy 50% + Hard 50% 0.7895 0.8276 0.7652
Easy 75% + Hard 75% 0.7866 0.8303 0.7580
Easy 100% + Hard 100% 0.7952 0.8302 0.7731
Table 2: Fine-tuning Results for Cellpose, StarDist, and CellViT Models. This table consists of three subtables summarizing the instance segmentation performance metrics (F1 score, Precision, and Recall) for each foundation model across three data enrichment experimental settings (using only “easy” images, only pathologist-corrected “hard” images, and a combination of both). Each sub-table compares the baseline performance (highlighted in blue) of the model against its fine-tuned counterparts. The highest values of each evaluation metric for the fine-tuned models are highlighted in red.
Figure 6: Comparison of F1-score across training/annotation strategies. The horizontal axis represents the percentage (%) of labels included under different strategies. Cellpose achieves optimal performance with “easy” labels, StarDist with “hard” labels, and CellViT with both. In this context, “hard” cases refer to training instances where automatic segmentation fails human quality assurance (QA) and employ pixel-level corrections, whereas “easy” cases are those where automatic results pass human QA.